Let x = Integral (squ root (sec(t))^4 -1) (dt) from -pi/4 to pi/4
1
Expert's answer
2011-10-14T09:09:28-0400
x=Integral ((squ root (sec(t))^4 -1) (dt)) from -pi/4 to pi/4 = Integral(1/(cos t)^2-1)from -pi/4 to pi/4= tan(t) | -pi/4 pi/4 - pi/4
+pi/4=1+1=2